Show parallel forces with the same direction F1 = 200 N and F2 = 50 N. The distance between their points of application is l = 50cm. Determine the value and point of application for the resulting force.


1
Expert's answer
2021-03-08T08:31:09-0500
"F=F_1+F_2=200+50=250\ N"

"200x=50(50-x)\\x=10\ cm"